Tests were conducted using a knife-edged polycarbonate sheet added to the bottom of the intake tonight in order to improve the effectiveness of the intake system. New rails have been integrated into the shooter system. The tests are still underway and there have not been any conclusive results yet.
Due to issues with establishing connection to the robot students replaced the old cortex and troubleshooted faulty wiring. There is still work to be done and students will continue to work to make Shockwave functional again.
Invitations to the Silicon Valley Regional were sent to almost all of our sponsors tonight.
Pit members worked on creating an easier means to set-up and transport the LED strips used for Pit lighting. Their fragility and difficulty to set up significantly complicates setting up the pit Thursday morning.